OpenStudy (anonymous):

given that : f(3)=1 f '(3)=-5 g(3)=-8 g '(3)=9 calculate (f*g)'(3) and (f/g)'(3)

OpenStudy (anonymous):

\[(fg)'=f'g+g'f \] so fill in the numbers

OpenStudy (anonymous):

\[(fg)'(3)=-5\times -8+1\times 9\]

OpenStudy (anonymous):

i get -31

OpenStudy (anonymous):

49 for the first one, which is correct.

OpenStudy (anonymous):

right because i can't do artithmetic. 49 is right

OpenStudy (anonymous):

\[(\frac{f}{g})'=\frac{gf'-fg'}{g^2}\]

OpenStudy (anonymous):

got both, thanks :]

OpenStudy (anonymous):

so we get \[\frac{-8\times-5-9\times 1}{(-8)^2}\]

OpenStudy (anonymous):

i will leave arithmetic to you yw